The Wholesome Buffet offers an all-you-can-eat buffet meal for $33 per person. The restaurant employs ten salaried employees. Rent for the building, employee salaries, and other fixed costs for the restaurant are $63,500 per month when the restaurant serves meals up to 13,900 guests in that month. If the restaurant were to serve more than 13,900 guests in any given month, it would need to hire two additional servers and two additional kitchen staff for an additional fixed cost of $18,000 per month. The variable cost of food is $16 per guest. Read the requirements. Requirement 1. What is the restaurant's cost equation for estimating costs if the number of guests to be served is up to 13,900 guests per month? 33 x+$63500 Requirement 2a. If the restaurant were to serve 12,700 guests in May, what would be the estimated total cost (Use the cost equation from Requirement If the restaurant were to serve 12,700 guests in May, the estimated total cost would be 482600 Requirement 2b. If the restaurant were to serve 12,700 guests in May, what would be the average cost per guest? If the restaurant were to serve 12,700 guests in May, the average cost per guest would be $ Requirement 3a. If the restaurant were to serve 10,700 guests in June, what would be the estimated total cost using the cost equation from 38 Requirement 1? If the restaurant were to serve 10,700 guests in June, the estimated total cost using the cost equation from Requirement 1 would be$ 416600 Requirement 3b. If the restaurant were to serve 10,700 guests in June, what would be the estimated total cost using the average per guest cost from Requirement 2b? If the restaurant were to serve 10,700 guests in June, the estimated total cost using the average cost per guest from Requirement 2b would be S 406600
Requirement 3b. If the restaurant were to serve 10,700 guests in June, what would be the estimated total cost using the average per guest cost from Requirement 2b? f the restaurant were to serve 10,700 guests in June, the estimated total cost using the average cost per guest from Requirement 2b would be $ Requirement 3c. Why is there a difference in the two cost estimates? There is a difference in the two cost estimates because the average cost per guest remains constant as volume changes, due to the fixed component of the buffet's costs. The fixed cost per unit remains constant as volume decreases, while the variable cost per unit decreases 406600 Requirement 4. What is the restaurant's cost equation for estimating costs if the number of guests to be served is more than 13,900 guests per month? у-3163500|x+$ 13900 Requirement 5. If the restaurant were to serve 15,300 guests in July, what would be the estimated total cost? (Use the cost equation from Requirement 4 If the restaurant were to serve 15,300 guests in July, the estimated total cost using the cost equation from Requirement 4 would be$
Answer #1

1) y= 16x + $ 63,500

2)Total cost = 12,700*16 +63,500= 266,700

2b) Average cost per guest =266,700/12,700 = 21

3a)Total cost = 10,700*16 +63,500 = 234,700

3b)Total cost =10,700*21 = 224,700

3c)There is a Difference in two cost estimates because the average cost per guest changes as volume changes , due to the Fixed Component of the buffets cost. The fixed Cost per unit increases as volume decreases. while the variable cost per unit Remains constant.

4) y= 16x + 81,500

5)Total cost = 15,300*16 +81,500 =326,300